# A Guide to Misinformation Detection Datasets

### v1

@@ 5.1 Baseline Performance
We next discuss baselines when using these datasets. Recent works have shown LLMs represent the state-of-the-art for misinformation detection on generic claims Chen & Shu (2023); Pelrine et al. (2023). But despite its importance Pelrine et al. (2021), both human and compute time constraints can be a barrier to comparing with such strong baselines.
Thus, we provide two baselines for future use. We follow the recent method of Tian et al. (2024) and use GPT-4-0125 in two ways: directly prompting the LLM for a veracity evaluation, and providing the LLM a web search tool to first collect evidence before forming a final verdict. We note that although these are state-of-the-art systems for zero-shot misinformation detection, they should not be regarded as sole or permanent points of comparison. Stronger LLMs and methods could replace them eventually. Nonetheless, they can provide a useful point of comparison for the near future.
@@
5.2 The Flaw in Current Metrics
When looking at the outputs of the prediction system, we observe cases where the predicted label did not match the ground truth, yet the evidence and reasoning of the system was valid. For instance, in one example on the FEVER dataset, the input claim is “Vietnam is a place” and the prediction said roughly “Vietnam is not just a place, it’s a country!” In another example from LIAR-New, a statement was marked false by PolitiFact because it was in the context of a fake video, but the statement itself did not mention the video and in isolation would be true. In cases like these (and further examples in Appendix A.6), a simple binary or categorical label cannot provide an informative evaluation.
To determine the prevalence of this phenomenon, two annotators manually evaluated (Appendix A.7) chain-of-thought rationales from the web-search enabled baseline prediction system Tian et al. (2024).
Table 6: Agreed-upon manual annotations. Many predictions marked invalid by simple comparison with the categorical labels are actually valid.
@@ 5.2 The Flaw in Current Metrics
Rationale is Valid Rationale is Invalid Rationale is Valid Rationale is Invalid
@@
We observed a consistently high false-incorrect rate (first column of Table 6) and a generally low false-correct rate (fourth column of Table 6). Therefore, when benchmarking generative AI misinformation detection systems using categorical labels, the predictive accuracy and similar metrics reflect a reasonable lower bound on the performance—but a terrible upper one. We also observe that there is a large amount of ambiguity and room for interpretation in the examples that are being marked wrong by categorical label in these three datasets. Hence:
  1. 1.
    Categorical metrics cannot be used alone to compare generative and non-generative systems. Although multiple recent works (e.g., Pelrine et al. (2023); Chen & Shu (2023); Wei et al. (2024); Yu et al. (2023)) have highlighted the effectiveness of recent LLMs for misinformation detection, their comparisons with prior approaches may even still be underestimating the dominance of LLMs in this domain.
  2. 2.
@@
Although to our knowledge not addressed in the context of misinformation detection on claims, challenges in evaluation of generative AI have been broadly documented in other fields (McIntosh et al., 2024; Ahuja et al., 2023; Michel-Villarreal et al., 2023; Basole & Major, 2024). A common approach aimed at solving this is having an LLM evaluate the full prediction rationales (Bai et al., 2022; Sun et al., 2023b), with Sun et al. (2023a) noting that LLM-powered evaluation can produce more consistent preference signals than human annotators. In general, using LLMs for evaluation enables one to leverage much richer signals than simple categorical predictions and labels, while avoiding reliance on often inaccessible human evaluators.
@@
As an initial step towards higher-fidelity evaluation, we constructed an evaluator based on contradictions between the explanation generated by a predictive system, and a fact-checking article. In particular, we provide GPT-4-0409 both the prediction and the fact-checking article, and ask it to score contradictions from 0 (no contradiction) to 10. The exact prompt and other implementation details are provided in Appendix A.8. We chose a score-based approach to avoid forcing a potentially misleading binary in cases where there is a partial contradiction. With this approach, good predictions should have low contradiction against a high quality, professional fact-checking article. We also tested binary and trinary versions of this prompt, described in the Appendix, which yielded nearly identical results.
@@
To evaluate this evaluator, we set the oracle-optimal threshold of 3 or less (low contradiction) to indicate a prediction that the rationale is not wrong, and 4 or more indicating one that the rationale is wrong. The contradiction score evaluator then agrees 68% of the time with the human labels of valid and invalid rationales (reported in Table ) on the LIAR-New dataset. This is higher than the 60% original human interannotator-agreement on these rationales (before disagreement resolution described in Appendix A.7), and suggests the method extracts a meaningful but not definitive evaluation signal. We also note, though, that there is more to high quality misinformation detection than just a lack of contradiction. Therefore, we do not suggest using this tool as a primary evaluator. But we provide these results, along with all the data—inputs, predictions, manual labels, fact-checking articles, and evaluator outputs—on our GitHub, as a potential springboard for stronger evaluation methods in future research.
6 Limitations
@@
First, we note that while to our knowledge this represents the largest and most comprehensive survey of datasets in this domain, there are certainly many other datasets in existence and it is probable that some were not included. There is also a steady stream of new datasets being created. In the near future, we plan to collect external feedback and update our survey to maintain and expand the comprehensiveness of our study.
@@
Lastly, although we discussed multiple key aspects of misinformation detection datasets, there are still more that are worth considering. For example, much of our analysis focuses on claims datasets, which are unquestionably important but by no means the entirety of valuable data in the field. Similarly, although dataset quality in terms of labels and spurious correlations is critical, there can be other important considerations like ambiguous statements (Pelrine et al., 2023). We aim to address some of these in future work.
7 Conclusion
@@
High quality data is essential for realistic results and rapid progress in the field. In this work, we have provided a guide to misinformation detection datasets aiming at both quantity and quality. It also highlights limitations of existing datasets and evaluation approaches, which may have uncertain labels, spurious correlations, and misleading results. We hope that on the one hand this work can provide a roadmap for future methods research that needs to select datasets and evaluation approaches, and on the other, provide the foundational understanding and call to action to improve the misinformation detection dataset landscape.
Acknowledgments
This work was partially funded by the CIFAR AI Chairs Program, the Centre for the Study of Democratic Citizenship, the Social Sciences and Humanities Research Council of Canada (grant no. 435-2023-0628) and the Digital Citizen Contribution Program of Canadian Heritage. Kellin Pelrine was supported by funding from IVADO and by the Fonds de recherche du Québec.
Author Contributions
Camille Thibault led the dataset collection and curation, and the keyword analysis, and contributed to the predictive output annotation. Gabrielle Peloquin-Skulski led the label analysis and unification, and the partisanship and ideological lean analysis. Jacob-Junqi Tian led the baselines and the engineering for many of the experiments. All three made major contributions to the writing and other parts of the project. Florence Laflamme did both annotation and case studies of predictive outputs. Yuxiang Guan built the project website and other key contributions to making the data accessible. Reihaneh Rabbany and Jean-François Godbout advised the project. Jean-François Godbout also contributed to the writing and project management. Kellin Pelrine advised and managed the project, led the temporal correlations analysis, and proposed many of the key research directions including the project itself and that evaluation with categorical label metrics might be systemically flawed.
Appendix A Appendix
A.1 Supplement on Data Collection
@@
A.2 Dataset Details
This section provides an overview of datasets containing claims. The number of claims, the collection method, and the original labels are discussed.

### v2

@@ 4.1. Baseline Performance
We next discuss baselines when using these datasets. Recent works have shown LLMs represent the state-of-the-art for misinformation detection on generic claims (Chen and Shu, 2023; Pelrine et al., 2023). But despite its importance (Pelrine et al., 2021), both human and compute time constraints can be a barrier to comparing with such strong baselines.
Thus, we provide two new baselines for future use. We follow the recent method of (Tian et al., 2024) and use GPT-4-0125 in two ways: directly prompting the LLM for a veracity evaluation, and providing the LLM a web search tool to first collect evidence before forming a final verdict. We note that although these are state-of-the-art systems for zero-shot misinformation detection, they should not be regarded as sole or permanent points of comparison. Stronger LLMs and methods could replace them eventually. Nonetheless, they can provide a useful point of comparison for the near future.
@@
4.2. The Flaw in Current Metrics
When looking at the outputs of the prediction system, we observe cases where the predicted label did not match the ground truth, yet the evidence and reasoning of the system was valid. For instance, in one example on the FEVER dataset, the input claim is “Vietnam is a place” and the prediction said roughly “Vietnam is not just a place, it’s a country!” In another example from LIAR-New, a statement was marked false by PolitiFact because it was in the context of a fake video, but the statement itself did not mention the video and in isolation would be true. In cases like these (and further examples in Appendix C.2), a simple binary or categorical label cannot provide an informative evaluation.
To determine the prevalence of this phenomenon, two human expert annotators evaluated (Appendix C.3) chain-of-thought rationales from the web-search enabled baseline prediction system (Tian et al., 2024).
Table 3. Agreed-upon manual annotations. Many predictions marked invalid by simple comparison with the categorical labels are actually valid. Standard evaluations are likely to systematically under-evaluate the performance of generative systems.
@@ 4.2. The Flaw in Current Metrics
Rationale is Valid Invalid Valid Invalid
@@
We observe a consistently high false-incorrect rate (first column of Table 3) and a generally low false-correct rate (fourth column of Table 3). Therefore, when benchmarking generative AI misinformation detection systems using categorical labels, the predictive accuracy and similar metrics reflect a reasonable lower bound on the performance—but a terrible upper one, that marks many valid responses incorrect. We also observe that there is a large amount of ambiguity and room for interpretation in the examples that are being marked wrong by categorical label in these three datasets. Hence:
  1. (1)
    Categorical metrics cannot be used alone to compare generative and non-generative systems. Although multiple recent works (e.g., Pelrine et al. (2023); Chen and Shu (2023); Wei et al. (2024); Yu et al. (2023)) have highlighted the effectiveness of recent LLMs for misinformation detection, their comparisons with prior approaches may even still be underestimating the dominance of LLMs in this domain.
  2. (2)
@@
Although to our knowledge not addressed in the context of misinformation detection on claims, challenges in evaluations of generative AI have been broadly documented in other fields (McIntosh et al., 2024; Ahuja et al., 2023; Michel-Villarreal et al., 2023; Basole and Major, 2024). A common approach aimed at solving this is having an LLM assess the full prediction rationales (Bai et al., 2022; Sun et al., 2023b), with Sun et al. (2023a) noting that LLM-powered evaluation can produce more consistent preference signals than human annotators. In general, using LLMs for evaluation enables one to leverage much richer signals than simple categorical predictions and labels, while avoiding reliance on often inaccessible human evaluators.
@@
As an initial step towards higher-fidelity evaluation, we constructed an evaluator based on contradictions between the explanation generated by a predictive system, and a fact-checking article. In particular, we provide GPT-4-0409 both the prediction and the fact-checking article, and ask it to score contradictions from 0 (no contradiction) to 10. The exact prompt and other implementation details are provided in Appendix C.4. We chose a score-based approach to avoid forcing a potentially misleading binary in cases where there is a partial contradiction. With this approach, good predictions should have low contradiction against a high quality, professional fact-checking article. We also tested binary and trinary versions of this prompt, described in the Appendix, which yielded nearly identical results.
@@
To evaluate this evaluator, we set the oracle-optimal threshold of 3 or less (low contradiction) to indicate a prediction that the rationale is not wrong, and 4 or more indicating one that the rationale is wrong. We find that the contradiction score evaluator agrees 68% of the time with the human labels of valid and invalid rationales (reported in Table 3) on the LIAR-New dataset. This is higher than the 60% original human inter-annotator agreement on these rationales (before disagreement resolution described in Appendix C.3), and suggests the method extracts a meaningful but not definitive evaluation signal. We also note, though, that there is more to high quality misinformation detection than just a lack of contradiction. Therefore, we do not suggest using this tool as a primary evaluator. But we provide these results, along with all the data—inputs, predictions, manual labels, fact-checking articles, and evaluator outputs—on our GitHub, as a potential springboard for stronger evaluation methods in future research.
5. Evaluation Quality Assessment
We have highlighted multiple ways datasets and evaluation procedures could deliver results that will fail to translate to the real world. While the information here directly reveals some pitfalls, we suspect that future research—with new datasets, new methods, and new needs—will encounter similar problems. To improve the validity of conclusions in our field in a systematic and future-proof way, we propose the Evaluation Quality Assessment (EQA) as a key component of an improved research methodology.
We draw inspiration from the common practices for systematically and transparently documenting dataset publications, such as “datasheets for datasets” (Gebru et al., 2021), and suggest that similar attention should be given to the usage of datasets. However, it is important to ensure data and evaluation quality not just as a lengthy post-hoc appendix, but as an integral part of the research process and a key foundation of every paper’s core arguments. Therefore, we propose including EQA as a scientifically mandatory step within the methodology. This means augmenting the standard descriptive discussion of data used with a critical analysis of it, explaining what steps and quality assessments were conducted to assure that the data is suitable to prove the experimental conclusions. We suggest that dedicating even a paragraph to this, rather than taking quality of “standard” datasets for granted, could dramatically improve the validity of our empirical results and the rate of real progress in the field. Conversely, current practices risk misleading evaluations, generalization failures, and immeasurable time and effort dedicated to measurements of noise.
We make three specific recommendations for Evaluation Quality Assessment as follows.
  • Papers should express clearly the bounds on the intended generalization of their method, and in turn what types of generalization the evaluation aims to prove. For example, at one extreme, if the evaluation data is a perfect IID sample of the application data and there is no distribution shift whatsoever, then many spurious correlations and other quality issues are irrelevant. But most methods are intended to apply to data with distributional differences in time, topic, source, or other characteristics. By briefly but concretely laying out which distributional shifts are in scope and which are not in scope, a paper states their evaluation hypotheses.
  • Papers should report analyses performed on the data and evaluation process to assure it is suitable to test the intended generalization of the method. One universal baseline is that author should qualitatively inspect at minimum 25 random examples—input, output, and ground truth—where their method made correct predictions, and 25 where it made incorrect ones. This data should be provided for reproducibility, and provides a basic sanity check. We note, for example, that the feasibility and categorical metrics issues we have reported would be detected by such a sanity check in the majority of datasets we examined. Furthermore, papers should also perform quantitive assessments of their data and evaluation, such as (but not limited to) the ones discussed in prior sections and available in our CDL-DQA toolkit.
  • Papers should report limitations in EQA either directly in methodology or in the standard limitations section. Authors should consider not just surface-level limitations like the number or size of datasets in the experiments, but also potential evaluation quality issues that the EQA could not rule out. For instance, if generalization under temporal distribution shifts is expected but the method was trained on data not checked for spurious temporal correlations, this would represent a limitation that future research should be aware of and consider addressing.
In this way, we treat assessing evaluation as a crucial step for validating the research process. For authors, this will help reinforce the demonstrated rigour of the experiments and the validity of proposed methods. It will also prevent the rejection of strong methods due to unrealistically high baseline settings caused by spurious correlations or other evaluation issues. For reviewers, this will produce more explicit grounding (or lack thereof) for the experiments, as well as a framework to recommend when evaluation is insufficiently justified and needs better validation. Thus, we propose that adopting EQA can efficiently and substantively mitigate propagating and compounding errors arising in this domain from current practices of copy-pasting evaluation from past literature.
6. Related Work
In recent years, the scientific community has devoted considerable research to curbing misinformation, defined broadly as “false or misleading information” (Lazer et al., 2018). Initial surveys, such as Shu et al. (2017), provided foundational frameworks and overviews of detection techniques from data mining perspectives, while subsequent studies by Oshikawa et al. (2018) and Zhou and Zafarani (2020) emphasized NLP methods and interdisciplinary approaches. Despite these contributions, most existing surveys (Bondielli and Marcelloni, 2019; Gravanis et al., 2019) largely overlook detailed analyses of misinformation datasets, offering superficial coverage that neglects dataset-specific qualities essential for accurate model development and validation. Recent surveys partially address this gap by categorizing datasets quantitatively (D’Ulizia et al., 2021; Sharma et al., 2019; Ali et al., 2022; Patra et al., 2022), yet typically lack in-depth analysis. Other works, including Abdali (2024), Pelrine et al. (2021), and Wu and Hooi (2022), provide deeper insights into dataset biases and limitations but remain limited by analyzing only a small subset of datasets. To address these limitations, our work evaluated a broad selection of misinformation datasets, assessed in depth how issues in data and evaluation could lead to misleading conclusions, and provided tools and recommendations for future improvements in the field.
7. Limitations
@@
First, we note that while to our knowledge this represents the largest and most comprehensive survey of datasets in this domain, there are certainly many other datasets in existence and it is probable that some were not included. There is also a steady stream of new datasets being created every year. In the near future, we plan to collect external feedback and update our survey to maintain and expand the comprehensiveness of our study.
@@
Lastly, although we discussed multiple key dimensions of misinformation detection datasets, a favorable assessment in these dimensions does not guarantee a dataset suitable for any application. Other types of limitations in the data could still lead to spurious shortcuts instead of generalizable predictions and evaluation. Our proposed EQA procedure can help detect this with respect to particular methods and applications, but we also encourage future work to identify additional limiting factors in evaluation in this domain, and to solve them.
8. Conclusion
@@
High quality data and evaluation are essential for realistic results and rapid progress in the field. In this work, we have provided a guide to misinformation detection datasets aiming at both quantity and quality. We also highlighted limitations of existing datasets and evaluation approaches, which may have spurious correlations, infeasible examples, and misleading results. We hope that this work can provide a roadmap to better grounding for future predictive methods research that needs to select datasets and evaluation approaches. Meanwhile, we also hope that this guide will provide foundational understanding and a call to action to build new and better datasets and evaluation procedures.
Appendix A Survey and Construction of CDL-MD
In this section, we provide an expanded discussion of our collection of combined datasets on true and false information, including a total of 75 datasets, with 36 specifically focused on claims and statements, and 9 specifically focused on paragraphs. The full dataset contains a total of 120,901,495 observations, while the subset that we further analyze includes 1,741,146 observations. These data encompass a wide range of topics, including political issues, health concerns, and environmental questions, often related to the United States but also covering international news, headlines and online posts. The original labels within the datasets were assigned through a combination of expert evaluations and algorithmic methods. The following section provides a detailed summary of the data collection process and the characteristics of these datasets.
A.1. Expanded Related Work
In recent years, the scientific community has shown a growing interest in fake news detection to mitigate the spread of misinformation, defined as “false or misleading information” (Lazer et al., 2018). Within this evolving field, several surveys have emerged to offer comprehensive reviews and standardized evaluations. A pioneering effort by Shu et al. (2017) provided an early framework, defining fake news, detailing its characteristics, and summarizing detection techniques from a data mining perspective. Subsequent surveys, such as Oshikawa et al. (2018) and Zhou and Zafarani (2020), have explored alternative methodologies, focusing respectively on natural language processing (NLP) methods and interdisciplinary perspectives. However, while these surveys and others (Bondielli and Marcelloni, 2019; Gravanis et al., 2019) are valuable for providing a comprehensive overview of the state-of-art in fake news detection, they pay limited attention to existing datasets. Indeed, even if some emphasize the challenges of data collection or stress the importance of dataset quality, these surveys usually provide only superficial coverage of existing datasets, overlooking their specific content, details, and characteristics. Assessing the quality of these misinformation datasets is critical because they are often used to train and test models for misinformation detection and related tasks. A lack of quality data in this context implies that biases and erroneous conclusions could be introduced both in the development stage and in the validation process of these systems.
This gap has thus spurred the emergence of additional surveys dedicated to addressing these dataset-centric nuances, which can be categorized into two types. The first one focuses on categorizing existing datasets to guide the research community in their selection. For example, D’Ulizia et al. (2021) surveyed 27 datasets based on eleven characteristics (e.g., application purpose, type of misinformation, language, size, news content type, etc.) and compared these quantitatively. Another example, Sharma et al. (2019) summarized the characteristic features of 23 existing datasets, providing a clearer picture of those available to the public. However, these surveys have an important drawback; they often lack in-depth analysis. In fact, only descriptive characteristics are listed, thus neglecting key characteristics of their quality and effectiveness for future research. This is also the case for Ali et al. (2022) and Patra et al. (2022), which describe 26 and 7 datasets, respectively.
The second type of survey focuses on analyzing the quality, performance, and limitations of datasets. For instance, Abdali (2024) examines 10 datasets to identify some of these weaknesses and strengths. However, a broad approach is used to outline biases in this study, which fails to detail the specifics of each dataset, leaving researchers uncertain about their individual quality. Another example, Hamed et al. (2023) highlight the limitations of 20 articles using publicly available datasets. While this approach provides a good overview of literature trends, a grey area remains regarding whether the errors in these 20 articles stem primarily from methodology or dataset issues. We also find the work of Pelrine et al. (2021), who evaluate the quality of six datasets, focusing on their potential spurious correlations with temporal information. Wu and Hooi (2022) expands the analysis of the spurious correlations issue to those induced by event-based collection, dataset merges, and labeling bias, using the Twitter15, Twitter16, and PHEME datasets. However, in both of these studies, the limited number of datasets analyzed fails to provide a comprehensive view of the diverse landscape of available datasets in this field. Similarly, Pelrine et al. (2023) highlight issues of ambiguous claims in the LIAR dataset, but does not expand their analysis beyond that one and their own LIAR-New dataset.
A.2. Collection Process
We then expanded our selection by rigorously examining the citations in scientific articles related to these initial sources. Some articles listed available datasets in their literature reviews or surveys, enabling us to incorporate additional data. Through these combined approaches—which are further detailed in the Appendix—we identified 75 publicly accessible datasets, presented in Table 4.
For our analyses, we refined our selection to focus on claims datasets and paragraph datasets. Claims datasets contain textual claims, defined here as short statements ranging from one to two sentences. Tweets are included in this definition, while lengthier online and social media posts are excluded. Paragraph datasets contain more comprehensive arguments, defined by having more than two sentences. Articles and social media posts fit under this definition. Our selection excludes hybrid datasets, which are defined as possessing data types which can be classified as both claims and paragraphs. These two dataset types are qualitatively different because statements and claims are more concise than paragraphs, such as OP-ED and news articles, which often include opinions, commentary, and contextual details. This extraneous information can potentially obscure the core claim or statement and introduce noise in the labeling process, as information can be partly true or false.
A.3. Supplement on Data Collection
@@
A.4. Claims Datasets
A summary description of each of the claims datasets can be found in Appendix A.7.1. Of these datasets, 12 consist of claims scraped from fact-checking or reliable websites, another 12 consist of tweets, and the remaining 11 comprise claims drawn from Twitter, the internet, social media, or news websites. There is variation in the topics of these datasets, but most focus on areas with significant societal impact where misinformation is prevalent and potentially harmful. For example, 16 of the datasets focus on health, vaccination, and COVID-19; 3 focus on political issues; 1 on environmental issues, and the rest covers various subjects, from culture, sport, the economy and so on. Unfortunately, a significant limitation of much of this data is the absence of information regarding the date the claim was made or fact-checked. This can potentially impact the accuracy of labeling, given that certain claims may have been true or false at the time they were made. In addition, this limitation affects the scope of our temporal leakage analysis. Consequently, scholars, and practitioners alike should be cautious when using these data.
A.5. Paragraph Datasets
A summary description of each of the paragraph datasets can be found in Appendix A.7.2. Among the 9 datasets, 8 of the datasets focused on general political misinformation and 1 of the datasets focused on pop culture misinformation. A notable difference in the labeling approaches between claims and paragraphs is that paragraphs tend to rely more on crowd-sourced and source-based labeling approaches, whereas claims datasets tend to rely more heavily on human experts. This makes sense to the extent that relying on human experts is unfeasible for large quantities of text. Paragraph datasets are defined as datasets that strictly contain data types that are either articles, Facebook posts, information from Wikipedia, news articles, or magazine articles. Unfortunately, a significant limitation is that there are many datasets where there is a mixture of claim data types and paragraph data types. There are 39 paragraphs that include paragraph data types, but only 9 of them purely contain paragraph data types without any claim data types. Since the other 30 hybrid datasets give no method of distinguishing between claim data types and paragraph data types in the datasets themselves, we exclude these hybrid cases from the set of paragraph datasets. This illustrates that there is a need for more pure paragraph datasets in the literature.
A.6. Labeling Approach
Harmonizing labels across different datasets is a crucial step to ensure comparability of results and robustness of analyses. Since each study employs its own criteria for classifying veracity (see Appendix A.7.1), in this survey, we use the original labels from the studies to create a more consistent categorical variable across datasets. Specifically, we classify content as true, false, mixed or unknown. Information that is mostly true is classified as true, while content that is mostly false is classified as false. Finally, ambiguous claims, such as those partially true or false, were classified as mixed, while claims that were unproven, unrelated or contained no information about their veracity were coded as unknown. The percent of true and false claims in each dataset using this coding scheme is shown in Table 5. Moreover, the original annotation method for each dataset, along with its advantages and disadvantages, is detailed in Appendix A.8.
A.7. Dataset Details
This section provides an overview of all claims datasets and paragraph datasets. The number of entries, the collection method, and the original labels are discussed.
A.7.1. Claims Dataset Details